Opened 10 years ago

Closed 4 years ago

Last modified 4 years ago

#331 closed flight processing (fixed)

FW10/02, flight day 101/2010, Cumbria

Reported by: benj Owned by: knpa
Priority: alpha 4 medium Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2010/flight_data/uk/FW10_02-2010_101_full_waveform_test

Data arrived from ARSF via network transfer on 12th April 2010.

Scientific objective: Test flight for full-waveform LiDAR.

Priority: a4m (not graded by SC, assigned following consultation with ppu)

PI: Doreen Boyd

Full waveform test flight. Will be delivered to PI but not specifically requested.

Sensors:

  • Eagle (19/08/2010)
  • Hawk (20/12/2010)
  • Leica LIDAR (Full Waveform) (06/08/2010)

Change History (42)

comment:1 Changed 10 years ago by mggr

  • Summary changed from FW10_02, flight day 101/2010, Cumbria to FW10/02, flight day 101/2010, Cumbria

Fix summary line (in case we ever automate this).

comment:2 Changed 10 years ago by mark1

IPAS data has been processed.

comment:3 Changed 10 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

comment:4 Changed 9 years ago by emca

No sync message for Hawk lines 20, 21 or 22

comment:5 Changed 9 years ago by emca

Per Line SCT values. GPT values, where applicable, are given second:

Line No: Eagle Hawk
01 -0.18, -0.18 +0.02
02 +0.07 +0.07
03 -0.14, -0.16 +0.04
04 +0.06 +0.04
05 -0.06 0
06 +0.06 +0.04
07 -0.10 +0.01
08 +0.05 +0.04
09 +0.01 -0.02
10 +0.02 +0.05
11 +0.01 +0.01
12 +0.05 +0.05
13 +0.04 -0.03
14 +0.04 +0.04
15 +0.06 0
16 -0.01, -0.12 +0.03
17 +0.08 -0.03
18 -0.09, -0.12 +0.03
19 +0.06 -0.03
20 -0.13, -0.12 +0.26
21 +0.03 +0.43
22 -0.12, -0.16 +0.90, +0.04
23 +0.08, +0.04 -0.03
24 -0.18, -0.18 +0.02
25 +0.19, +0.15 -0.01

comment:6 Changed 9 years ago by emca

The following photographs have been removed as they are outside the project area:
1-13, 135-138, 218, 219

comment:7 Changed 9 years ago by emca

Hyperspectral, Camera and LIDAR are ready for delivery check.

H: ~airborne/workspace/FW10_02-2010_101_Cumbria/delivery/20100706
C: ~airborne/workspace/FW10_02-2010_101_Cumbria/delivery/20100707
L: ~airborne/workspace/FW10_02-2010_101_Cumbria/delivery/20100708

comment:8 Changed 9 years ago by knpa

Beginning LiDAR delivery check.

comment:9 follow-up: Changed 9 years ago by knpa

Delivery checked LIDAR.

Looks good apart from the following points:

Logsheet has incorrect format, should be: FW10_02-2010_101_Cumbria.pdf
Logsheet shouldn't be executable.
Project code in Readme should be FW10_02
line 132 "These points have been classified rather than removed incase they are"
Why SRTM and not nextmap used for DEM?
Strange files in N1 and N2: os and sys.

Note: could not check LAS 1.3 Files as no .trj files in delivery. Should these be included?

comment:10 in reply to: ↑ 9 Changed 9 years ago by mggr

Replying to knpa:

Note: could not check LAS 1.3 Files as no .trj files in delivery. Should these be included?

Tricky.. it's specific to TerraScan but given that's the only tool we have, I'd agree they should be included.

Perhaps we should also include the sol/sbet file to give people another option for navigation info? Thoughts?

Note also that we should point out the LAS 1.3 files are not compliant to the specification, though are compatible with TerraScan - we'll have to redeliver these when Leica give us an updated processor.

comment:11 Changed 9 years ago by jb362

Beginning hyperspectral delivery check

comment:12 Changed 9 years ago by jb362

Hyperspectral delivery check complete.

1) Read_Me is in .txt format and should be in .pdf.

2) Logsheet should be named FW10_02-2010_101_Cumbria.pdf

Apart from that all fine.

Hawk band 160 blank for all flight lines.

comment:13 Changed 9 years ago by jb362

Beginning camera delivery check

comment:14 Changed 9 years ago by jb362

Camera delivery check complete. Everything fine.

comment:15 Changed 9 years ago by emca

LiDAR and Photography data dispatched to both D. Boyd and K. Anderson 06/08/2010

comment:16 Changed 9 years ago by emca

LiDAR and Photography have been rsync'd to
/users/rsg/arsf/arsf_data/2010/flight_data/uk/FW10_02-2010_101_Cumbria

Did not rsync the leica/proc_laser folder as the .LAS files were included in the delivery.

comment:17 Changed 9 years ago by emca

  • Description modified (diff)

Eagle data dispatched to PI 19/08/2010

comment:18 Changed 9 years ago by anch

  • Description modified (diff)

I have created FW extractions for some interesting areas in the flightlines including roads, trees, buildings etc.

There are inside delivery directory 20100924.

comment:19 Changed 9 years ago by emca

The corrected LAS1.3 files made available to D Boyd and K Anderson via ftp 21/10/2010

(incorrect header information on the original delivered LAS1.3 files due to a problem with the processing software)

comment:20 Changed 9 years ago by mark1

Have checked the Hawk delivery and all looks good. Just needs the bad pixel file updated when it is available.

comment:21 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data sent to PI 08/12/2010

comment:22 Changed 9 years ago by emca

Reprocessed the Hawk Lev1 files and added them to the 20101206 delivery folder

comment:23 Changed 9 years ago by mark1

Delivery checked and hawk data looks good, ready to deliver.

comment:24 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data re-delivered 20/12/2010

comment:25 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa

This project is ready for archiving but not completely r-synced. Need to clear out some space in repository first, I think around 30G needed.

comment:26 Changed 9 years ago by emca

Eagle & Hawk data sent to K. Anderson (via Jamie) on a NTFS formatted disk.
Also included a LiDAR dem as requested.

comment:27 Changed 9 years ago by emca

Received an email from D.Boyd to say there was an error with the HDF file for Eagle line 21.

This line has been reprocessed and the HDF file made available to D Boyd and K Anderson via FTP. Have replaced the bad file in the delivery on the workspace and in the repository.


comment:28 Changed 8 years ago by knpa

Remained of r-sync completed. Ready to be archived.

comment:29 Changed 8 years ago by knpa

  • Description modified (diff)

comment:30 Changed 8 years ago by knpa

  • Description modified (diff)

comment:31 Changed 7 years ago by knpa

Made available to NEODC 09/07/12

comment:32 Changed 7 years ago by emca

Reprocessing hyperspectral data using APL

No sync for Hawk line 21 and 22. New sct's for Hawk data:
line 20: 0.01
line 21: 0.42
line 22: 0.89

Processing all bands for Hawk data

comment:33 Changed 7 years ago by emca

Hawk line 9 had no hdr file with the updated wavelengths - created this

comment:34 Changed 7 years ago by emca

Hawk files created with no problems.

Have had to use gpt offsets to get some of the Eagle lines to match up (using -posattoff in aplnav)
New sct + gpt values for Eagle data
1) -0.22, +0.16
3) -0.24, +0.18
16) -0.16, +0.12
18) -0.15, +0.13
20) -0.22, +0.17
22) -0.20, +0.18
23) +0.03, -0.04
24) -0.25, +0.19
25) +0.16, -0.18

Reprocessing the Eagle data for all bands.

comment:35 Changed 7 years ago by emca

Updated sct values for two more hawk lines:

line 06: -0.07
line 13: -0.06

comment:36 Changed 7 years ago by emca

Created delivery for reprocessed hyperspectral data - ready for delivery check

comment:37 Changed 7 years ago by besm

Delivery checked, no problems found.

comment:38 Changed 7 years ago by anhi

Redelivery not requested, this is finished

comment:39 Changed 7 years ago by knpa

Original deliveries ingested by NEODC.

Moved to archive disk.

comment:40 Changed 7 years ago by knpa

Removed workspace version (original processing).

Leaving ticket open until reprocessing is archived.

comment:41 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Archived data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/FW10_02/FW10_02-2010_101_Cumbria). LiDAR data is duplicated - have notified Wendy.

Marking as closed.

comment:42 Changed 4 years ago by dac

Archiving

Duplicate LAS directory at NEODC removed.

Note: See TracTickets for help on using tickets.